OpenStudy (anonymous):

log 3 0.1 + 2log 3 x = log 3 2 + log 3 5??

OpenStudy (jiteshmeghwal9):

\[\log_{3}{0.1}+2\log_{3}{x}=\log_{3}{2}+\log_{3}{5}\]\[\log_{3}{1\over10}+2\log_{3}{x}=\log_{3}{2}+\log_{3}{5}\]now use the property\[\log_{a}{x}+\log_{a}{b}=\log_{a}{xb}\]\[\log_{3}{{1\over10}+x^2}=\log_{3}{2 \times 5}\]\[\implies \frac{1}{10}+x^2=6\]solve this expression & gt ur answer:)